9. An analysis of company performance using DuPont analysis Aa Aa Walking down the hall of your office building with a sheaf of papers in his hand, your friend and colleague, Akira, stepped into your office and asked the following. Akira: Do you have 10 or 15 minutes that you can spare? You: Sure, I've got a meeting in an hour, but I don't want to start something new and then be interrupted by the meeting, s So how can I help? performance, in Akira: I've been reviewing the company's financial statements and looking for general ways to improve our general, and the company's return equity, or ROE, in particular. Emma, my new team leader, suggested that I start by on C using a DuPont analysis, and I'd like to run my numbers and conclusions by you, to see if I've missed anything. the balance sheet and income statement data that Emma gave me, and here are my notes with my calculations. Here are Could you start by making sure that my numbers are correct? You: Give me a minute to look at these financial statements and to remember what I know about the DuPont analysis. Balance Sheet Data Income Statement Data Accounts payable Cash $800,000 $960,000 Sales $16,000,000 Accounts receivable Accruals 320,000 Cost of goods sold 1,600,000 8,000,000 Notes payable Gross profit Inventory 2,400,000 1,280,000 8,000,000 Current liabilities Operating expenses Current assets 4,800,000 2,560,000 4,000,000 Long-term debt 3,680,000 ЕBIT 4,000,000 Total liabilities 6,240,000 Interest expense 595,200 Common stock 840,000 ЕВT 3,404,800 Net fixed assets 4,800,000 Retained earnings 2,520,000 Taxes 1,191,680 Total equity $2,213,120 3,360,000 Net income Total assets Total debt and equity $9,600,000 $9,600,000
If I remember correctly, the DuPont equation breaks down our return on equity (ROE) into three component ratios: the net profit margin the total asset turnover ratio, and the equity multiplier And, according to my understanding of the DuPont equation and its calculation of ROE, the three ratios provide insights into the company's effectiveness in using the company's assets, and Now, let's see your notes with your ratios, and then we can talk about possible strategies that will improve the ratios. I'm going to check the box to the side of your calculated value if your calculation is correct and leave it unchecked if your calculation is incorrect. Cepeus Manufacturing Inc. DuPont Analysis Check if Check if Ratios Value Ratios Value Correct Correct Profitability ratios Asset management ratio Gross profit margin (%) 50.00 Total asset turnover 1.67 Operating profit margin (%) 21.28 Net profit margin (%) Financing ratios 23.05 equity (%) Equity multiplier Return on 59.28 1.54 Akira: OK, it looks like I've got a couple of incorrect values, so show me your calculations, and then we can talk strategies for improvement You: I've just made rough calculations, so let me complete this table by inputting the components of each ratio and its value:
You: I've just made rough calculations, so let me complete this table by inputting the components of each ratio and its value: Cepeus Manufacturing Inc. DuPont Analysis Ratios Calculation Value Profitability ratios Numerator Denominator Gross profit margin (%) Operating profit margin (%) Net profit margin (%) Return on equity (%) Asset management ratio Total asset turnover Financing ratios Equity multiplier Akira: I see what I did wrong in my computations. Thanks for reviewing these calculations with me. You saved me from a lot of embarrassment! Emma would have been very disappointed in me if I had hershowed my original work. So, now let's switch topics and identify general strategies that could be used to positively affect Cepeus's ROE. You: OK, so given your knowledge of the component ratios used in the DuPont equation, which of the following strategies should improve the company's ROE? (Check all that apply.) Increase the cost and amount of assets necessary to generate each dollar of sales because it will increase the company's total asset turnover Decrease the amount of debt financing used by the company, which will decrease the total asset turnover ratio. Increase the firm's bottom-line profitability for the same volume of sales, which will increase the company's net profit margin. Decrease the company's use of debt capital because it will decrease the equity multiplier.
Answer #1

1. The DuPont analysis breaks ROE into three component ratios: the net profit margin, the asset turnover ratio and equity multiplier

2. The Three ratios provides insights into the company's Shareholder and Dividend Management, Effectiveness in using and control over its expenses

3. Gross profit margin = Gross profit/Sales = 8000000 / 16000000 = 50%

4. Operating profit margin = Operating profit/Sales = 4000000 / 16000000 = 25%

5. Net profit margin = Net profit/Sales = 2213120 / 16000000 = 13.83%

6. Return on equity = Net profit/Total Equity = 2213120 / 3360000 = 65.87%

7. Total asset turnover = Sales/Total Assets = 16000000 / 9600000 = 1.67

8. Equity Multiplier = Total Assets/Total Equity = 9600000 / 3360000 = 2.86

The points that should improve ROE are -

Option D Decrease the company's use of debt capital

Option C. Increase the firm's bottom line profitability